OpenStudy (anonymous):

A ball is kicked upward with an initial velocity of 56 feet per second. The ball's height h (in feet) can be expressed as a function of time t (in seconds) by the equation h = -16t2 + 56t. How much time does the ball take to return to the ground?

OpenStudy (anonymous):

h(t) = 56 t - 16 t ² 0 = 56 t - 16 t² 16 t ² - 56 t = 0 (4 t) (4 t - 14 ) = 0 t = 0 , t = 3.5 roots will be start of kick till touch down